HELLO,
I AM STRUGGLING WITH LINEAR PROBLEMS. I DON'T KNOW WHICH SIDE I AM TO SUBTRACT. EXAMPLE: 6X-3Y=9 OR X-4Y=8 OR 4Y=9X-7. THANK YOU, I AM TAKING THIS ALGEBRA A SECOND TIME AND I REALLY WANT TO UNDERSTAND AND PASS THE CLASS.

Answer by mananth(16946)

I cant take a class in this forum.
But it is very simple . This is just a method and you have to follow them to solve the equations.
1. Bring all the x & y terms to one side of the equation.
Example : x = 2y +9
rewritten as x-2y =9 ( whenever a term or number changes it sides of the equation the sign changes.Here 2y has changed sides so it has become -2y)
2. The second equation is y=x+6.
rewritten y-x =6
Now solve these two equations for x & y